OpenStudy (anonymous):

it concerns the outflow of water from a cylindrical tank with a hole at the bottom you are asked to find the height of the water in the tank at any time if the tank has a diameter 2M, the hole has diameter 1CM and the initial height of the water when the hole is opened is 2.25M, when will the tank be empty???????

OpenStudy (ganpat):

Volume of water = pi * r2 * h = 28.274 meter cube Let x be amount of time taken to empty a tank, so per minute = 1/x Amount of water out, Q = pi / 4 * d2 * sq rt 2gh = = 3.14 /4 * 1/100 sq * sq (2 * 9.81 * 2.25) = 5.21 * (1/100) square.. this is per minute.. so total time = 28.274 * 100 * 100 / 5.21 =

OpenStudy (ganpat):

m not sure man.. just a try, also recheck on calculations..

OpenStudy (anonymous):

it's gd i'll check the answer

OpenStudy (ganpat):

= 5.21 * (1/100) square.. this is per minute.. This should be per sec.. FLow is mostly meter cube / sec... So whatever answers come divide by 60.. to get into minutes..

OpenStudy (anonymous):

thx ganpat